OpenStudy (anonymous):

Need help with quadratic equation: x*squared* - 5 = 2x

OpenStudy (anonymous):

do you know how to solve quadratic equation?

OpenStudy (anonymous):

basically what you want to do is put all the terms onto one side. therefore: \[x ^{2}-2x-5=0\] after that, what you want to do is identify the numbers in front of the terms [called coefficients]. infront of \[x ^{2}\] there's a 1. in front of x, there's a -2, and -5. so, A=1 B=-2 C=-5 therefore using the quadratic equation: \[\left( -b \pm \sqrt{b^{2}-4(a)(c)} \right) / (2)(a) \] plug in those values and solve. so: \[\left( 2 \pm \sqrt{4-4(1)(-5)} \right) / (2)(1) \] \[\left( 2 \pm \sqrt{24} \right) / (2) \] and so after simplifiying: \[\left( 2 \pm (2)\sqrt{6} \right) / (2) \] you can factor out a 2 and you'll also be canceling out a the denominator. \[\left( 2 (1\pm \sqrt{6} \right) / (2) \] the 2's cancel out leaving you with the answer: \[\left( 1\pm \sqrt{6} \right) \] so your answers are: \[\left( x=1+ \sqrt{6} \right) \] \[\left( x=1- \sqrt{6} \right) \]

OpenStudy (anonymous):

heh sorry that took a while, i type slow. : \